find the least common denominator of the rational expression with denomonators of 9x-63 and 12x-84.
i do no know how to find the correct way to do this.
thanks!
i hope you can help me! :)

Answer by stanbon(75887)

LCD of 9x-63 and 12x-84.
Factor each denominator as follows:
3^2(x-7) and 3*2^2(x-7)
The LCD is the product of the highest power of each of these prime factors:
LCD=3^2 * 2^2 * (x-7)= 36(x-7)
Cheers,
Stan H.